OpenStudy (anonymous):

complete explanation of cos45cos60-sin45sin60

OpenStudy (cwrw238):

this is identical to cos(45 + 60)

OpenStudy (cwrw238):

compound angle identity cos (A + B) = cosAcosB - sinAsinB

OpenStudy (anonymous):

how to evaluate this quistion

OpenStudy (cwrw238):

ok - its = cos 105 degrees using my calculator this = -0.2588

OpenStudy (cwrw238):

need any more help?

OpenStudy (anonymous):

cos45=1/sqrt2and cos60=1/2so i am putting their values (1/sqrt2*1/2-1/sqrt2*sqrt3/2) i have problem in their multiplication

OpenStudy (anonymous):

in next step it will be 1/2sqrt2 *sqrt3/2sqrt2 now explain how to take LCM in this case please

OpenStudy (cwrw238):

1/2sqrt2 - sqrt3/2sqrt2 = sqrt3/4

OpenStudy (cwrw238):

sorry the last part is wrong its (1 - sqrt3) / 2sqrt2

OpenStudy (cwrw238):

-this checks out ok against - 0.2588 which i quoted earlier

OpenStudy (anonymous):

but if you solve it on calculator its showing -sqrt6+sqrt2/2

OpenStudy (cwrw238):

no - there an error somewhere -sqrt6 + sqrt2/2 = -1.7424

OpenStudy (anonymous):

so explain me how you take LCM 2sqrt2 it can be 8

OpenStudy (cwrw238):